OpenStudy (anonymous):

need helping solving diff eq. dy/dx + y^1/3 = 0 y(0) = Yo

ganeshie8 (ganeshie8):

separate variables ?

OpenStudy (anonymous):

So I subtracted y^1/3 from both sides then seperated..

OpenStudy (anonymous):

dy/y^1/3 = -1dx

ganeshie8 (ganeshie8):

looks good, integrate

OpenStudy (anonymous):

this is where I am having some problems.......do I just move y up top as y^-1/3....this is where I am differeing from my classmates.

ganeshie8 (ganeshie8):

\[\frac{dy}{y^\frac{1}{3}} = -1 dx\] is same as \[y^\frac{-1}{3} dy = -1 dx\]

OpenStudy (anonymous):

so the integral on left side would be -1/3*y^2/3 right? My classmates are smehow getting 3/2y^2/3

ganeshie8 (ganeshie8):

\[\large \int y^\frac{-1}{3} dy = \int -1 dx\] \[\large \dfrac{y^{\frac{-1}{3}+1} }{\frac{-1}{3}+1} = -x + c\] \[\large \dfrac{y^{\frac{2}{3}} }{\frac{2}{3}} = -x + c\] \[\large \frac{3}{2}y^{\frac{2}{3}}= -x + c\]

OpenStudy (anonymous):

Thanks.....its been a long day...obviously lol

ganeshie8 (ganeshie8):

:) you still need to plugin the initial value and solve \(c\)

OpenStudy (anonymous):

i got it (;